If The Ksp Of AgBr Is 5.0 X 10-13, What Is The Molar Solubility (Mol/L) Of AgBr?

1
Expert's answer
2022-04-25T17:04:10-0400

Solution:

The Ksp of AgBr is 5.0×10−13


AgBr(s) ⇌ Ag+(aq) + Br(aq)

The Ksp expression for AgBr(s) is:

Ksp = [Ag+][Br]


ICE Table:



Substitute the equilibrium concentration values from the ICE Table into the Ksp expression: 

Ksp = [Ag+] × [Br] = (x) × (x) = x2 = 5.0×10−13

To find the molar solubility of AgBr, solve the equation for x:

x = (5.0×10−13)1/2 = 7.07×10−7

Molar solubility of AgBr = 7.07×10−7 M


Answer: The molar solubility of AgBr is 7.07×10−7 mol/L